# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Bridges

### v1

@@ 3.1 Identifying Bridge Transactions
Pairing these component transactions (i.e., matching deposits on one chain to withdrawals on other) can be performed in several different ways. The easiest, and most common, is via a unique transaction identifier. Such ids are typically generated by the bridge on the source chain during a deposit transaction and then copied into the withdrawal transaction on the destination chain. [^5] These unique ids are commonly simple global variables incremented with each new transaction. However, instead of an explicit id, some bridges use a hash of the deposit transaction for the same purpose (e.g., for Anyswap, each withdraw transaction will include the hash of its corresponding deposit transaction). [^6] We note that to prevent potential replay attacks, bridge implementers should also include an id that unique identifies individual deposit events within a transactions. Sadly, this is often not the case. Finally, in a handful of cases there are no “inband” identifiers that can be used to associate transactions. We believe this is a poor design choice that is fundamentally in conflict with transparency. However, even in these cases, for the purpose of our analysis we have been able to pair transactions using explicit query APIs provided by the affected bridges services. [^7] For example, when a bridge transaction on the Poly Network bridge includes a withdrawal from Curve (a kind of liquidity pool) or when a bridge transaction on the Binance Token Hub includes a Binance Smart Chain (BSC) withdrawal, it is not possible to identify the partner deposit from blockchain data alone and we must make use of “out of band” data available through their respective bridge query APIs.
@@
At the end of this process, we identify a comprehensive collection of “bridge transactions” (a pair of transactions from two different blockchains that were used by the bridge to transfer of value across them). While the vast majority of bridge transactions are pairs of deposit and withdrawal transactions, a handful of bridge transactions are either withdrawal-only (e.g., in the case of attacks) or deposit-only (e.g., if the user chooses to delay their withdrawal).
3.2 Identifying the Value Transferred
@@
Many bridge transaction event formats explicitly identify the amount and kind of tokens transferred. However, some bridges do not emit this information and others can be unreliable. [^8] For example, as mentioned earlier, pre-2024 Wormhole does not emit an event for withdrawal transactions. Other examples include the Meter bridge and Qubit bridge, which attempted to verify the number of tokens received but had exploitable bugs. In such cases, we can frequently make use of the ERC-20 “Transfer” event (see Figure 3 for an example) that is emitted when the contract transfers the tokens from the user’s account to the bridge’s account. This event contains the number of tokens transferred as well as the sender and recipient addresses, which we use to identify the number of tokens transferred to or from the bridge. Because the Transfer event is adjacent to its associated bridge-generated event, it is easy to identify and thus establish the number of tokens transferred. [^9] As a sanity check, we also require that the sender and recipient addresses in the Transfer event are consistent with the bridge’s defined behavior: for withdrawal transactions, we require that the sender is either a mint address (i.e., all-zero address) or a bridge-controlled address, while for deposit transactions, we require that the recipient is either a burn address (i.e., all-zero address) or a bridge-controlled address. In a few implementations, multiple related Transfer events can be emitted at once, and in these cases we have manually inspected their contracts and constructed implementation-specific logic to account for this behavior. Another special case is caused by so-called “reflection tokens” in which the number of tokens logged in the event (the value field in Figure 3) is dynamically adjusted based on a combination of the intended number and the total token supply. For such cases, we either rely on bridge events which capture the intended number of tokens transferred or implement token-specific logic to recompute the value accordingly. Finally, native tokens have no Transfer event (since they are not ERC-20 tokens), but thus far we have either been able to recover the number of tokens transferred from bridge events or so-called “internal transactions”.
@@
To summarize, while it is certainly possible to create a bridge transaction protocol that records insufficient data to match deposits and withdrawals, or for which the number of tokens transferred might be ambiguous, our empirical experience analyzing 11 bridges and 20 blockchains is that such reconstruction has always been possible.
3.3 Checking the Balance Invariant
@@
Table 1: The blockchains CrossChecked supports and the cross-chain bridges that operate on them. While a particular attack on a bridge involves two chains, we collect deposit and withdrawal transactions for a bridge on all chains that the bridge supports.
@@
Ronin Mar 2022 $624.0$m 3.0m 2 - - - -
PolyNetwork/2021 Aug 2021 $611.0$m 292k 18 - - 1 -
BSC Token Hub Oct 2022 $587.0$m 2.0m 2 - - - -
Wormhole Jan 2022 $360.0$m 642k 1 - - 2 -
Nomad Aug 2022 $152.0$m 37k 962 - - - -
Harmony Jun 2022 $100.0$m 336k 15 - - - 43
HECO Nov 2023 $86.0$m 23k 8 - - - 73
Qubit Jan 2022 $80.0$m 260 16 - 114 - -
Anyswap Jul 2021 $7.9$m 3.4m 4 24 - 806 7
PolyNetwork/2023 Jun 2023 $4.4$m 290k 136 - - 1 27
Chainswap Jul 2021 $4.4$m 53k 1136 - 15 4 -
Meter Feb 2022 $4.3$m 14k 5 - - - -
Total $2.6$b 10.1m 2,305 24 129 814 150
@@
Table 2: List of top attacks on cross-chain bridges in the retrospective analysis, ordered by amount stolen. CrossChecked analyzed over 10m bridge transactions (20m component deposit and withdrawal transactions) and identified all bridge transactions previously reported as having been associated with the attacks (Reported). It also identified bridge transactions that violated the invariant that were a previously unidentified attack (New), test transactions (Test), transactions reflecting bugs in implementations or use (Error), and suspicious transactions that employ manual signing (Suspicious). CrossChecked identified slightly more transactions than were reported by the Nomad [20] and Meter [22] bridges for their attacks (+2 for Nomad, +1 for Meter). The Chainswap attack only has reports of the malicious deposit address [18], which matches the one identified by CrossChecked.
4 Retrospective Analysis
@@
The basic reasoning motivating the balance invariant is simple: that legitimate bridge transactions should conserve value. However, this assumes a particular model for how bridges are used and operated which may or may not hold in practice. To validate our hypothesis, we applied the balance invariant analysis retrospectively across a large set of past transactions which, while primarily benign, contain the largest known bridge attacks during our period of study. Our goal is both to show that all real attacks are identified (detection), but also that the invariant does not alert on large numbers of benign transactions (bridge compatibility). Later, in Sections 5 and 6, we will describe systems for live bridge monitoring as a third party and a new implementation for preventing unbalanced transactions from being committed.
4.1 Data Set
@@
To perform a retrospective analysis we need to data for which (at least some) of the ground truth is known. In this section we describe how we chose the bridge attacks we study, the blockchains and smart contracts involved, and how we collected the historical transaction data.
Attack Selection.
@@
We compiled a comprehensive list of attacks on cross-chain bridges that occurred between January 2021 and December 2023. These were sourced both from academic papers surveying cross-chain bridge attacks [14, 36, 38] as well as industry blog posts that collect and characterize attacks over time [1, 27, 6, 32, 31]. As this paper focuses on end-to-end auditing of bridge transactions, we filter out attacks involving blockchains lacking smart contracts such as Bitcoin (e.g., the pNetwork attack in 2021), attacks on swap bridges (e.g., the Thor bridge attacks), attacks that do not involve individual bridge transactions (e.g., the evoDefi attack), and attacks that withdraw funds through other means (e.g., direct transfer from a vault that stores assets for a bridge such as the Multichain 2023 attack). Moreover, given the large number of attacks during this period, we focus on major attacks with claimed losses greater than \$1 million USD and exclude the remaining (the sum of losses from these excluded attacks are a small fraction of those in our scope). This process yielded 12 attacks on 11 distinct bridges. We note that this list includes the top five attacks on cross-chain bridges in history, which collectively resulted in more than \$2.6 billion USD in claimed losses.
Blockchain Selection.
@@
Validating bridge transactions requires access to transaction data from both the source and destination chains. We support every blockchain involved in the bridge transactions associated with the 12 attacks, either as the source (or claimed source) or destination chain. As a result, we support a total of 21 blockchains that together cover a broad range of designs, including many of the most popular such as Ethereum, Binance Smart Chain, and Solana. [^10] Details about some deposit transactions on PolyNetwork (deposits that are used for liquidity purposes) are only available through PolyNetwork’s API. We support querying PolyNetwork’s API, but do not list it as a supported blockchain. Table 1 lists the blockchains and the bridges in our retrospective study that operate on them.
Smart Contract Selection.
@@
For each bridge, we comprehensively collected the results of all versions of its bridging smart contracts on every blockchain we considered. In particular, we collected deposit and withdrawal transactions created by every contract the bridge deployed on every blockchain we support (typically many more blockchains than the ones involved in a bridge attack) as well as all versions of the smart contract implementation for the bridge (bridges evolve their implementations over time, such as in response to an attack). Since the transactions we collected from this set of smart contracts are much broader than those just involved in the top 12 attacks we consider, including them further reinforces our findings that the alerting workload is very small (Section 4.2).
@@
Figure 5: The lifetime of the bridges in our retrospective study. Lines start with the bridge’s first valid transaction and end with the last valid transaction in our data, corresponding to the bridge’s closure or November 2023 if the bridge was still operating at the end of our data set. Diamonds indicate the dates of attack.
Data Collection.
@@
Collecting smart contract transaction data—the verbose record of what each contract did—across many chains constitutes the most time-intensive aspect of performing a retrospective analysis.
@@
For each bridge, we collected deposit and withdrawal transactions for each of the blockchains it operates on (for the 21 chains we support) using commercial RPC services which charge for queries. We primarily used five commercial services, as no single service supports all of the blockchains and functionality we need. [^11] These five RPC services are: GetBlock, QuickNode, ChainStack, GoldRush, and Ankr. For each bridge and blockchain combination, by default we collected all of the historical transactions generated by its smart contracts from the genesis of its deployment to the end of November 2023 or the end of its life, whichever came first.
@@
The two exceptions to this rule are the Binance bridge, for which we limited data collection to a year (six months prior to its attack and six months after) because of the sheer volume of transactions over its lifetime (30 million transactions) and the Harmony bridge, whose contract was reused by another bridge (LayerZero) after its attack (since LayerZero was not attacked itself, we only consider transactions that were directly associated with Harmony).
@@
4.2 Analysis
We developed a tool called CrossChecked that pairs deposit and withdrawal transactions from blockchains into bridge transactions, and applies the balance invariant and consistency checks on them. Using the transactions we collected for the 12 attacks across 11 bridges and 21 chains, CrossChecked analyzed over 10m bridge transactions, identifying more than 2.3k bridge transactions associated with the attacks and 1.1k more that violated the balance invariant.
Table 2 summarizes our results. For each attack, it shows the bridge involved, the date of the attack, the claimed loss in USD, the number of transactions CrossChecked analyzed, and the number and kinds of transactions that violated the balance invariant. Below we discuss these two categories of transactions in more detail. For further context on each of the attacks, Appendix A provides a descriptive account summarizing how the attacks transpired.
Looking forward to Sections 5 and 6, altogether CrossChecked identified 3,423 bridge transactions (0.03%) that violated the balance invariant out of more than 10m analyzed. If the invariant is used by a third-party auditing or protection system, we note that such an alert workload has a negligible overhead for manual inspection, typically raising no more than one alert (or one batch of alerts) every few weeks.
4.2.1 Reported
For each of the 12 significant cross-chain attacks, Section 4.1 describes how we gathered the historical transactions that correspond to the attacks using external sources. We use these identified transactions as ground truth for evaluating the ability of CrossChecked to identify attack transactions. As shown in Table 2, when processing the transaction histories of the chains involved, CrossChecked successfully identified all 2,305 bridge transactions on the source and destination chains associated with the attacks.
@@ 4.2.1 Reported
4.2.2 Other Violating Transactions
Equally compelling is the extent to which other, non-attack transactions violate the balance invariants. If the attack transactions are dominated by many false positives, then the approach becomes less effective. As shown in Table 2, CrossChecked finds significant fewer (1,117 compared to 2,305) bridge transactions that were not previously identified as attacks. Given the nature of these additional transactions, though, they do not undermine the effectiveness of the invariant approach. By violating the invariant something highly unusual is taking place. As a result, we argue that such transactions should be flagged for further scrutiny and perhaps even blocked from executing (particularly transactions in the “new” category and the large transactions in the “suspicious” category below).
To ensure that we have not missed benign explanations for a violation, we manually inspected violating bridge transactions in at least of one of the following ways: (1) using blockchain explorers to verify that the funds have been created by the withdrawal (and if so, whether they have been moved); (2) searching online for any additional information about the transaction and addresses involved; (3) checking that if claimed deposit exists; and (4) examining unredeemed deposits on the source chain that potentially could have been used to back the withdrawal (e.g., because the smart contract implementation changed between the deposit and withdrawal). If we can manually reconcile a bridge transaction, we consider it benign and do not consider it further.
We group the remaining bridge transactions that violate the invariant into four categories, which we describe below. For reference, we also list some of these bridge transactions in Table 3 in the appendix to provide specific examples with more detail.
New. We believe that we have identified previously unreported bridge transactions involved in two new, unreported attacks on Anyswap. The first group of three transactions executed once Anyswap reopened after the attack but before Anyswap patched its smart contract (Anyswap transactions #1–3 in Table 3 in the appendix). These transactions were three days later and involved different deposit and withdrawal addresses than the July 10, 2021 attack. The second group of 21 transactions on November 18, 2021, were all withdrawing on Avalanche. These transactions either referenced deposits that had already been redeemed months previously (Anyswap #6 in Table 3) or referenced non-existing deposits (Anyswap #7 in Table 3). Manually inspecting the smart contract used, the attackers were exploiting a bug in Anyswap’s implementation (the access control code was commented out and thus ineffective) and one of the addresses was labeled as “KyberSwap Exploiter”.

### v3

@@ 4.1. Identifying and PairingToken Bridge Transactions
Pairing these component transactions (i.e., matching deposits on one chain to withdrawals on other) can be performed in several different ways. The easiest, and most common, is via a unique transaction identifier. Such IDs are typically generated by the bridge on the source chain during a deposit transaction and then copied into the withdrawal transaction on the destination chain. [^5] These unique IDs are commonly simple global variables incremented with each new transaction. In addition, instead of an explicit ID, some bridges use a hash of the deposit transaction for the same purpose (e.g., for Anyswap, each withdraw transaction will include the hash of its corresponding deposit transaction). Finally, in a handful of cases there are no “inband” identifiers that can be used to associate transactions. We believe this is a poor design choice that is fundamentally in conflict with auditability. However, even in these cases, for the purpose of our analysis we have been able to pair transactions using explicit query APIs provided by the affected bridge’s services, which represents the best information available. [^6] For example, when a bridge transaction on the Poly Network bridge includes a withdrawal from Curve (a kind of liquidity pool) or when a bridge transaction on the Binance Token Hub includes a Binance Smart Chain (BSC) withdrawal, it is not possible to identify the partner deposit from blockchain data alone and we must make use of “out of band” data available through their respective bridge query APIs.
@@
At the end of this process, we identify a comprehensive collection of “bridge transactions”: a pair of transactions from two different blockchains that were used by the bridge to transfer value across them. While the vast majority of bridge transactions are pairs of deposit and withdrawal transactions, a handful of bridge transactions are either withdrawal-only (e.g., in the case of attacks) or deposit-only (e.g., if the user chooses to delay their withdrawal).
4.2. Identifying the Value Transferred
Prior work (Hu et al., 2024) that also studies cross-chain bridges relies on replaying transactions to automatically identify the amount of tokens transferred. However, this approach is not practical for our retrospective analysis, which spans 21 blockchains. For one, some blockchains (e.g., EVMOS) are poorly developed and do not offer easy-to-access replay APIs. Additionally, some transactions cannot be replayed as it is not supported by the blockchain (e.g., Binance Beacon Chain). We show in Section 6 that value extraction can be fully automated when replaying is possible on well-developed blockchains (e.g., Ethereum and BSC).
@@ 4.2. Identifying the Value Transferred
Instead, for our retrospective analysis, we rely on a combination of bridge events and ERC-20 token transfer events, which can be analyzed programmatically. In cases where a bridge’s transaction event is reliable and explicitly contains the amount and type of tokens transferred, we use that information directly. In cases where this is not true, we make use of the ERC-20 “Transfer” event (see Figure 3 for an example). This event contains the number of tokens transferred as well as the sender and recipient addresses, which we use to identify the number of tokens transferred to or from the bridge. While a Transfer event can also be unreliable in the case of malicious tokens (e.g., tokens that lie in the Transfer event), we have not encountered such cases in our analysis, as bridges typically do not allow arbitrary tokens. Because the Transfer event is adjacent to its associated bridge-generated event, it is easy to identify and thus establish the number of tokens transferred. [^7] As a sanity check, we also require that the sender and recipient addresses in the Transfer event are consistent with the bridge’s defined behavior: for withdrawal transactions, we require that the sender is either a mint address (i.e., all-zero address) or a bridge-controlled address, while for deposit transactions, we require that the recipient is either a burn address (i.e., all-zero address) or a bridge-controlled address. Another special case is caused by so-called “reflection tokens” in which the number of tokens logged in the event (the value field in Figure 3) is dynamically adjusted based on a combination of the intended number and the total token supply. For such cases, we either rely on bridge events which capture the intended number of tokens transferred or implement token-specific logic to recompute the value accordingly. Finally, native tokens have no Transfer event (since they are not ERC-20 tokens), but thus far we have either been able to recover the number of tokens transferred from bridge events or so-called “internal transactions”.
In a few implementations, multiple related Transfer events can be emitted at once, and in these cases we have manually inspected their contracts and constructed implementation-specific logic to account for this behavior. In total, we had to manually determine the amount of tokens transferred for a few hundred transactions. This manual work was necessary as we could not replay these transactions, and it reflects the limitations in historic data quality, not our methodology. We later show in Section 6 that such manual work can be avoided entirely when replaying is possible.
@@
To summarize, while it is certainly possible to create a bridge transaction protocol that records insufficient data to match deposits and withdrawals, or for which the number of tokens transferred might be ambiguous, our empirical experience analyzing 11 bridges and 21 blockchains is that such reconstruction has always been possible.
4.3. Checking the Balance Invariant
@@
Table 1. The blockchains CrossChecked supports and the cross-chain bridges that operate on them. While a particular attack on a bridge involves two chains, we collect deposit and withdrawal transactions for a bridge on all chains that the bridge supports. Anyswap is also known as Multichain.
@@
Ronin Mar 2022 624.0624.0m 3.0m 2 - - - -
PolyNetwork/2021 Aug 2021 611.0611.0m 292k 18 - - 1 -
BSC Token Hub Oct 2022 587.0587.0m 2.0m 2 - - - -
Wormhole Jan 2022 360.0360.0m 642k 1 - - 2 -
Nomad Aug 2022 152.0152.0m 37k 962 - - - -
Harmony Jun 2022 100.0100.0m 336k 15 - - - 43
HECO Nov 2023 86.086.0m 23k 8 - - - 73
Qubit Jan 2022 80.080.0m 260 16 - 114 - -
Anyswap Jul 2021 7.97.9m 3.4m 4 24 - 240 29
PolyNetwork/2023 Jun 2023 4.44.4m 290k 136 - - 1 27
Chainswap Jul 2021 4.44.4m 53k 1136 - 15 4 -
Meter Feb 2022 4.34.3m 14k 5 - - - -
Total 2.62.6b 10.1m 2,305 24 129 248 167
@@
Table 2. List of top attacks on cross-chain bridges in the retrospective analysis, ordered by amount stolen. CrossChecked analyzed over 10m bridge transactions (20m component deposit and withdrawal transactions) and identified all bridge transactions previously reported as having been associated with the attacks (Reported). It also identified bridge transactions that violated the invariant that were a previously unidentified attack (New), test transactions (Test), transactions reflecting bugs in implementations or use (Error), and suspicious transactions that employ manual signing (Suspicious). CrossChecked identified slightly more transactions than were reported by the Nomad (Nomad, ) and Meter (PANews, ) bridges for their attacks (+2 for Nomad, +1 for Meter). The Chainswap attack only has reports of the malicious deposit address (Liquid, ), which matches the one identified by CrossChecked.
5. Retrospective Analysis
@@
The basic reasoning motivating the balance invariant is simple: that legitimate bridge transactions should conserve value. However, this assumes a particular model for how bridges are used and operated which may or may not hold in practice. To validate our hypothesis, we applied the balance invariant analysis retrospectively across a large set of past transactions which, while primarily benign, contain the largest known bridge attacks during our period of study. Our goal is both to show that all real attacks are identified (detection), but also that the invariant does not alert on large numbers of benign transactions (bridge compatibility). Later, in Sections 6 and 7, we will describe systems for live bridge monitoring as a third party and a new implementation for preventing unbalanced transactions from being committed.
5.1. Data Set
@@
To perform a retrospective analysis we need data for which (at least some) of the ground truth is known. In this section we describe how we chose the attacks we study, the blockchains and smart contracts involved, and how we collected the historical transaction data.
@@
Attack Selection. We compiled a comprehensive list of attacks on cross-chain token bridges that occurred between January 2021 and December 2023. We used both academic papers surveying cross-chain bridge attacks (Lee et al., 2023; Zhang et al., 2023; Zhao et al., 2023) as well as industry blog posts that collect and characterize attacks over time (0xDatapunk, ; SlowMist, ; defi, ; White, ; Whinfrey, ). As described in our threat model (Section 3), we filter out attacks involving blockchains lacking smart contracts such as Bitcoin (e.g., the pNetwork attack in 2021 (pNetwork Team, 2024)), attacks on swap bridges (e.g., the Thor bridge attacks (THORChain, 2024a, b)), attacks that do not involve individual bridge transactions (e.g., the evoDefi attack (Carreras and Dalton, 2024)), and attacks that withdraw funds without a bridge component (e.g., direct transfer from a vault as in the Multichain 2023 attack (Multichain, 2024)). Moreover, given the large number of attacks during this period, we focus on major attacks with claimed losses greater than \$1 million USD and exclude the remaining (the sum of losses from these excluded attacks are a small fraction of those in our scope). This process yielded 12 attacks on 11 distinct bridges. We note that this list includes the top five attacks on cross-chain bridges in history, which collectively resulted in more than \$2.6 billion USD in claimed losses.
@@
Blockchain Selection. Validating bridge transactions requires access to transaction data from both the source and destination chains. We support every blockchain involved in the bridge transactions associated with the 12 attacks, either as the source (or claimed source) or destination chain. As a result, we support a total of 21 blockchains that together cover a broad range of designs, including many of the most popular such as Ethereum, Binance Smart Chain, and Solana. Table 1 lists the blockchains and the bridges in our retrospective study that operate on them.
@@
Smart Contract Selection. For each bridge, we comprehensively collected the results of all versions of its bridging smart contracts on every blockchain we considered. In particular, we collected deposit and withdrawal transactions created by every contract the bridge deployed on every blockchain we support (typically many more blockchains than the ones involved in a bridge attack) as well as all versions of the smart contract implementation for the bridge (bridges evolve their implementations over time, such as in response to an attack). Since the transactions we collected from this set of smart contracts are much broader than those just involved in the top 12 attacks we consider, including them further reinforces our findings that the alerting workload is very small (Section 5.2).
@@
Figure 4. The lifetime of the bridges in our retrospective study. Lines start with the bridge’s first valid transaction and end with the last valid transaction in our data, corresponding to the bridge’s closure or November 2023 if the bridge was still operating at the end of our data set. Diamonds indicate the dates of attack.
@@
Data Collection. Collecting smart contract transaction data—the verbose record of what each contract did—across many chains constitutes the most time-intensive aspect of performing a retrospective analysis.
@@
For each bridge, we collected deposit and withdrawal transactions for each of the blockchains it operates on (for the 21 chains we support) using commercial RPC services which charge for queries. We primarily used five commercial services (GetBlock, QuickNode, ChainStack, GoldRush, and Ankr.), [^8] GetBlock, QuickNode, ChainStack, GoldRush, and Ankr. as no single service supports all of the blockchains and functionality we need. For each bridge and blockchain combination, by default we collected all of the historical transactions generated by its smart contracts from the genesis of its deployment to the end of November 2023 or the end of its life, whichever came first.
@@
There are two exceptions to this rule. For the Binance bridge, we limited data collection to a year (six months prior to its attack and six months after) because of the sheer volume of transactions over its lifetime (30 million transactions) and the limited time available before the sunset of the Binance Beacon Chain. For the Harmony bridge, whose contract was reused by another bridge (LayerZero) after its attack (since LayerZero was not attacked itself, we only consider transactions that were part of Harmony).
@@
5.2. Analysis
We developed a tool called CrossChecked that pairs deposit and withdrawal transactions from blockchains into bridge transactions, and applies the balance invariant and consistency checks on them. [^9] As consistent with prior work (Hu et al., 2024), we ignore transactions for which we cannot pair because of incomplete information (e.g., deposits to a valid blockchain that we do not support). Using the transactions we collected for the 12 attacks across 11 bridges and 21 chains, CrossChecked analyzed over 10m bridge transactions (20m individual deposit and withdrawal transactions), identifying more than 2.3k bridge transactions associated with the attacks and 568 more that violated the balance invariant.
Table 2 summarizes our results. For each attack, it shows the bridge involved, the date of the attack, the claimed loss in USD, the number of transactions CrossChecked analyzed, and the number and kinds of transactions that violated the balance invariant. Below we discuss these two categories of transactions in more detail. For further context on each of the attacks, Appendix B provides a descriptive account summarizing how the attacks transpired.
Looking forward to Sections 6 and 7, altogether CrossChecked identified 2,873 bridge transactions (0.03%) that violated the balance invariant out of more than 10m bridge transactions analyzed. If the invariant is used by a third-party auditing or protection system, we note that such an alert workload has a negligible overhead for manual inspection, typically raising no more than one alert (or one batch of alerts) every few weeks.
5.2.1. Known
For each of the 12 significant cross-chain attacks, Section 5.1 describes how we gathered the historical transactions that correspond to the attacks using external sources. We use these identified transactions as ground truth for evaluating the ability of CrossChecked to identify attack transactions. As shown in Table 2, when processing the transaction histories of the chains involved, CrossChecked successfully identified all 2,305 bridge transactions on the source and destination chains associated with the attacks (including a few ones that are missed by all public reports).
@@ 5.2.1. Known
5.2.2. Other Violating Transactions
Equally compelling is the extent to which other, non-attack transactions violate the balance invariants. If the attack transactions are dominated by many false positives, then the approach becomes less effective. As shown in Table 2, CrossChecked finds significantly fewer (568 compared to 2,305) bridge transactions that were not previously identified as attacks. Given the nature of these additional transactions, though, they do not undermine the effectiveness of the invariant approach. By violating the invariant something highly unusual is taking place. As a result, we argue that such transactions should be flagged for further scrutiny and perhaps even blocked from executing (particularly transactions in the New category and the large transactions in the Suspicious category below).
To ensure that we have not missed benign explanations for a violation, we manually inspected violating bridge transactions in at least one of the following ways: (1) using blockchain explorers to verify that the funds have been created by the withdrawal (and if so, whether they have been moved); (2) searching online for any additional information about the transaction and addresses involved; (3) checking that if a claimed deposit exists; and (4) examining unredeemed deposits on the source chain that potentially could have been used to back the withdrawal (e.g., because the smart contract implementation changed between the deposit and withdrawal). If we can manually reconcile a bridge transaction, we consider it benign and do not consider it further.
We group the remaining bridge transactions that violate the invariant into four categories, which we describe below. For reference, we also list some of these bridge transactions in Table 3 in the appendix to provide more detail.
New. We believe that we have identified previously unreported bridge transactions involved in two new, unreported attacks on Anyswap. The first group of three transactions executed once Anyswap reopened after the attack but before Anyswap patched its smart contract (Anyswap transactions #1–3 in Table 3 in the appendix). These transactions were three days later and involved different deposit and withdrawal addresses than the July 10, 2021 attack (yet utilizing the same compromised key). The second group of 21 transactions on November 18, 2021, were all withdrawing on Avalanche. These transactions either referenced deposits that had already been redeemed previously (Anyswap #6 in Table 3) or referenced non-existing deposits (Anyswap #7 in Table 3). Manually inspecting the smart contract used, the attackers were exploiting a bug in Anyswap’s implementation (missing access control code) and one of the addresses was labeled as “KyberSwap Exploiter”. These results also match those reported by Hu et al. (Hu et al., 2024).